# Which Demographic Features Are Relevant for Individual Fairness Evaluation of U.S. Recidivism Risk Assessment Tools?

### v2

@@ Abstract.
conference: ; June 16–20, 2025; Chicago, ILccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Governmental regulations
@@ 1. Introduction
However, technical fairness criteria might be hard to enforce without legal incentives. Nguyen et al. (2025) reviewed primary sources of U.S. state/federal law and found that while three major technical fairness criteria (procedural, group, and individual fairness) can be traced back to judicial interpretation of the U.S. “Equal Protection” and “Due Process” constitutional clauses, individual fairness has yet been operationalized in specific statutes or regulations. To facilitate the development of more concrete legal criteria on individual fairness, we conduct an original human subjects experiment to determine whether individual fairness should consider or ignore demographic features like race, sex, or age in the individual similarity function (to identify pairs of highly similar individual, so that the outcomes within each pair of similar defendants can be compared to check if these outcomes are similar enough to be “fair”).111 We store Supplementary Materials (e.g., survey form, data, Python code) in a Google Drive folder: https://drive.google.com/drive/folders/1OHyRIQ21ECEZwD2i1xE_7Ahzhqt9w1IO?usp=drive_link
Our work contributes to the rich literature by the Human - Computer Interaction (HCI) community on humans’ perception of algorithmic fairness. For example, Wang et al. (2020) studied how different demographic groups of participants perceive fairness differently when an AI model makes demographically biased predictions in the participants’ favors, a phenomenon named as “outcome favorability bias”. Harrison et al. (2020) conducted survey-based experiments where participants were asked to compare two models representing trade-offs between fairness-related properties. They observed a marginal preference for equalizing false positives across demographic groups over equalizing accuracy. Saxena et al. (2019) explored people’s attitudes about individual fairness, investigating which definitions people perceive to be the fairest in the context of loan decisions, and whether fairness perceptions change with the addition of sensitive information (i.e., race of the loan applicants). Grgic-Hlaca et al. (2016) surveyed 100 Mechanical Turk workers’ fairness perception when using different features in an RRA task. These studies underscore the complexity of determining relevant fairness criteria when demographic factors are considered.
@@ 3.2. Main Experiment
Therefore, we conducted a between-subjects experiment on Qualtrics (with participants recruited from Prolific). To enhance ecological validity and data quality, we apply the following participant filters: based in the U.S., English proficiency, Prolific approval rate of at least $98\%$, and Bachelor’s degree. We also provide fairness voting bonuses to incentivize participants to align their fairness evaluation with common sense and optional rationale bonuses to incentivize them to think carefully about each fairness judgment. Furthermore, we filter out responses from participants who fail two or more attention checks. The median completion time among participants who passed our attention checks is $18.5$ minutes. We pay each participant \$$\$3.7$, corresponding to $12$ USD per hour.
@@ 4.2. Quantitative analysis: no significant fairness judgement differences across conditions
Table 1. The “Fair” rating rate per condition. Each of $N=234$ participants who pass all attention checks gives $14$ ratings.

### v3

@@ Abstract.
conference: ; June 16–20, 2025; Chicago, ILjournalyear: 2025copyright: ccconference: 20th International Conference on Artificial Intelligence and Law; June 16–20, 2025; Chicago, IL, USAbooktitle: 20th International Conference on Artificial Intelligence and Law (ICAIL 2025), June 16–20, 2025, Chicago, IL, USAdoi: 10.1145/3769126.3769199isbn: 979-8-4007-1939-4/25/06ccs: Human-centered computing Empirical studies in HCIccs: Social and professional topics Governmental regulations
@@ 1. Introduction
However, technical fairness criteria might be hard to enforce without legal incentives. Nguyen et al. (2025) reviewed primary sources of U.S. state/federal law and found that while three major technical fairness criteria (procedural, group, and individual fairness) can be traced back to judicial interpretation of the U.S. “Equal Protection” and “Due Process” constitutional clauses, individual fairness has yet been operationalized in specific statutes or regulations. To facilitate the development of more concrete legal criteria on individual fairness, we conduct an original human subjects experiment to determine whether individual fairness should consider demographic features like race, sex, or age in the individual similarity function. [^1] We store Supplementary Materials (e.g., survey form, data, Python code) in a Google Drive folder: https://drive.google.com/drive/folders/1OHyRIQ21ECEZwD2i1xE_7Ahzhqt9w1IO?usp=drive_link
Our work contributes to the rich literature on humans’ perception of algorithmic fairness. For example, Wang et al. (2020) studied how different demographic groups of participants perceive fairness differently when an AI model makes demographically biased predictions in the participants’ favors. Harrison et al. (2020) conducted surveys where participants were asked to compare two models representing trade-offs between fairness-related properties. They observed a marginal preference for equalizing false positives across demographic groups over equalizing accuracy. Saxena et al. (2019) explored people’s attitudes about individual fairness, investigating which definitions people perceive to be the fairest in the context of loan decisions, and whether fairness perceptions change with the addition of sensitive information (i.e., race of the loan applicants). Grgic-Hlaca et al. (2016) surveyed 100 Mechanical Turk workers’ fairness perception when using different features in an RRA task. These studies underscore the complexity of determining relevant fairness criteria when demographic factors are considered.
@@ 3.2. Main Experiment
Therefore, we conducted a between-subjects experiment on Qualtrics (with participants recruited from Prolific). To enhance ecological validity and data quality, we apply the following participant filters: based in the U.S., English proficiency, Prolific approval rate of at least $98\%$, and Bachelor’s degree. We also provide fairness voting bonuses to incentivize participants to align their fairness evaluation with common sense and optional rationale bonuses to incentivize them to think carefully about each fairness judgment. Furthermore, we filter out responses from participants who fail two or more attention checks. The median completion time among participants who passed our attention checks is $18.5$ minutes. We pay each participant $\mathdollar 3.7$, corresponding to $12$ USD per hour.
@@ 4.2. Quantitative analysis: no significant fairness judgement differences across conditions
Table 1. The “Fair” rating rate per condition. Each of $N=234$ participants who pass all attention checks gives 1414 ratings.